# Training the Untrainable: Introducing Inductive Bias via Representational Alignment

### v1

@@ Appendix F Test Accuracy across Training
Figure 7: Testing accuracy improves across guidance for sequence modeling. We visualize the testing accuracy for sequence modeling as an example to demonstrate that guidance improves accuracy across training and this improvement is significantly better across training. This allows for another intepretation of the method outside of cross-entropy loss.
@@
We plot accuracies over training as a complement to cross-entropy loss in fig. 7 for the sequence modeling experiments. We can use these experiments to de-couple our results from properties of cross-entropy loss that may lead to misleading improvements across training. We find that accuracies improve consitently across training, supporting the loss curve interpretation that guided training improves results.
Appendix G Mean-Squared Error Loss
Figure 8: Guidance improves training with MSE loss. We apply guidance to a new loss function, mean-squared error, over the CIFAR-10 dataset. We use the Deep FCN as our target network and an untrained ResNet-18 as our guide network. We find that the performance of the Deep FCN saturates but improves when using a guide. This shows that guidance is not dependent on the particular loss function used.
Throughout this work, we have mainly focused on a classification task and employed the cross-entropy loss function as our main task loss. However, other loss functions can be applied which do not have the same properties as cross-entropy loss for learning classification tasks as well as other kinds of tasks like regression tasks. One such loss is mean-squared error (MSE) loss. We include an additional experiment with guidance where we train the Deep FCN for image classification with a ResNet-18 guide.
We design an experiment where we train Deep FCN on the CIFAR-10 dataset (Krizhevsky et al., 2010) with an untrained ResNet-18 guide. Our reasoning for using CIFAR-10 instead of ImageNet stems from prior work which demonstrates difficulty with ImageNet training while using a vanilla MSE loss (Weng et al., 2022). We train the Deep FCN both with and without a guide for 100 epochs with a batch size of 256.
We show results in fig. 8. Similar to fig. 2, we find that guidance signficantly improves performance of the Deep FCN in comparison to base training. Most notably, the Deep FCN saturates on CIFAR-10 during training. However, with ResNet-18 as a guide, the loss continues to decrease across training. This demonstrates that guidance is useful for other loss functions and is unaffected by the task loss.
Appendix H Guidance with RSA
To further understand the generality of guidance as an approach, we include one experiment where we guide a network using a different metric for representational alignment. We choose representational similarity analysis (RSA) (Kriegeskorte et al., 2008) as our main approach. Our main criteria for a metric is differentiability and compute time. RSA is differentiable and can be computed quickly per batch. This makes it a natural candidate after CKA. We first describe RSA and use it as part of guidance between ResNet-18 and the Deep FCN.
H.1 Representational Similarity Analysis
@@
We use the RSA formulation as described in Kriegeskorte et al. (2008). Specifically, RSA constructs representational dissimilarity matrices (RDMs) for two sets of representations and compares them using an outer similarity function.
@@
As with CKA, we use the complement of the similarity to construct $\bar{\mathcal{M}}$. In practice, we define $d$ to be the cosine distance between every pair of inputs and $s$ to be the pearson correlation between the RDMs as done in previous work (Conwell et al., 2021a).
H.2 Results
Figure 9: Guidance with RSA as the representational similarity metric maintains similar performance to CKA. We include a furher experiment where we change the metric for representational alignment from CKA to RSA during guidance training. We apply this to the Deep FCN with ResNet-18 as a guide network. We see that, like CKA, RSA alignment also allows for transferring the prior from ResNet-18. However, unlike CKA, the untrained guide network only does marginally better than the trained network, potentially indicating the RSA is better at transferring trained features.
@@ H.2 Results
We apply guidance with RSA to Deep FCN as our target network and ResNet-18 as our guide network. Similar to our CKA results, we train for 100 epochs with a batch size of 256, as RSA is sensitive to the number of samples when comparing sets of representations. We see results over the training, validation, and representational dissimilarity loss in fig. 9. The Deep FCN guided by a trained ResNet-18 achieves a top-5 accuracy of 11.02% and the Deep FCN guided by a randomly initialized ResNet-18 achieves a top-5 accuracy of 11.74%.
@@
We can first observe that guided training improves over base training as noted in fig. 2 and table 2. This demonstrates the generality of our approach to other metrics. As long as a representational similarity metric is differentiable, we can optimize the metric for alignment between two networks as a method to transfer the prior of one network to another.

### v2

@@ Appendix G Test Accuracy across Training
Figure 11: Testing accuracy improves across guidance for sequence modeling. We visualize the testing accuracy for sequence modeling as an example to demonstrate that guidance improves accuracy across training and this improvement is significantly better across training. This allows for another interpretation of the method outside of cross-entropy loss.
@@
We plot accuracies over training as a complement to cross-entropy loss in fig.˜11 for the sequence modeling experiments. We can use these experiments to de-couple our results from properties of cross-entropy loss that may lead to misleading improvements across training. We find that accuracies improve consistently across training, supporting the loss curve interpretation that guided training improves results.
Appendix H Basic Distillation Comparison
Figure 12: Distillation does not prevent overfitting. We compare basic distillation [33] to see if we can prevent overfitting. We use Deep FCN as our student network. We find that distillation with a trained ResNet-18 teacher network leads to a small improvement in performance but still has some patterns of overfitting. Distillation with an untrained ResNet-18 teacher network hurts performance.
To show the effectiveness of guidance, we compare it with distillation from [33]. Distillation involves transferring knowledge from a performant teacher network to a less performant student network via maximizing the alignment of the output logits. This encourages the student to have similar predictions as the teacher network. This occurs via the following loss function. Assume $\bm{Q}$} is the logits extracted from the target (student) network and the $\bm{P}$} is the logits extracted from the guide (teacher) network.
$\mathcal{L}_{\text{distill}}=\alpha*T^{2}*\text{KL}(\sigma({\bm{Q}}/T)||\sigma({\bm{P}}/T))+(1-\alpha)*\mathcal{L}_{\text{CE}}({\bm{Q}},y)$ (8)
where $y$ is the ground truth labels, $T$ is the temperature to soften the logits, and $\alpha$ is the weighting factor between the distillation loss and cross-entropy loss. In this case, KL refers to the Kullback-Liebler divergence and $\sigma$ corresponds with the softmax function. In practice, we set $\alpha$ to 0.5 and $T$ to 2. We continue to track the full cross-entropy loss across training as well.
We show accuracy-based results in fig.˜5 and loss curve results in fig.˜12. Distillation with a trained network can improve performance but much less than guidance. Distillation with an untrained network reduces performance on average, although not by a significant amount.
H.1 Error Consistency
Figure 13: Distillation lowers error consistency. In general, we find that distillation results in error consistency patterns that are less consistent than what is reported with guidance.
We show error consistency performance over distilled networks rather than guided networks in fig.˜13.
Appendix I Guided Network Analysis and Interpretation
The results from guidance open many questions in order to explain why untrained guide networks can be better at improving target network performance. We provide an intuitive explanation as well as some geometric analysis of guided networks to see if there is a stronger interpretation.
I.1 Interpreting Guidance
We believe several prior works support our findings and interpretations in this paper. We cover them here. First, [38] is a recent paper that considers task-aware representational alignment. Their theory provides a generalization bound via kernel alignment. They show that when a “stitcher” maps representations or a source network to a target output, the excess risk of the stitched model is upper-bounded by the CKA alignment between them. This provides a learning-theoretic guarantee that the CKA term in guidance reduces the hypothesis class possibilities seen by an optimizer. Overfitting or underfitting becomes harder.
Shan and Bordelon [70] investigates how a network’s neural tangent kernel (NTK) aligns with a target output during training. The paper shows that NTK alignment accelerates convergence and lowers generalization error in deep linear networks. This aligned kernel condition is inserted by hand in guidance. Similarly, [3] uses Rademacher complexity tools to show that alignment of tangent-kernel features onto a small set of task-relevant directions compresses the effective model class. This formalizes the notion of guidance as an automatic regularizer, where task directions are replaced by the guide network settings. Finally, [37] demonstrates that after training, the top singular vectors of a network’s hidden activations align with the task target vectors. This empirically supports the layerwise CKA choice in guidance. We believe that CKA bounds the risk or complexity in terms of kernel alignment. The NTK and Rademacher analyses show that alignment shrinks the effective hypothesis space and improves conditioning. This aligns with findings based on singular vectors. We could sharpen this theory by changing the alignment used in guidance e.g. moving from aligning on kernels to aligning on singular vectors or eigenvectors instead. A full PAC-style proof specialized to guidance has not been shown in our paper but we leave this to future work.
I.2 Geometric Analysis via Intrinsic Dimensionality
Figure 14: Guidance preserves intrinsic dimensionality, avoiding over-regularization. We measure the PCA-based intrinsic dimensionality of representations from each layer of both a guided and unguided Deep FCN at initialization i.e. Random FCN, 300 steps of training, and the end of training. We find that guidance with an untrained guide network better preserves intrinsic dimensionality in comparison to base training or using a trained guide. All networks collapse to the same intrinsic dimensions. This establishes that guidance does change the dynamics of training based on geometric features of the target network.
We aim to understand how guidance with a randomly initialized guide network differs from a trained guide network. To do so, we compare the representation space of a target network guided by a trained guide network and a randomly initialized guide network using intrinsic dimensionality.
Intrinsic dimensionality (ID) refers to the minimum number of dimensions required to capture the structure or variability in the input data. It represents the true complexity of the data manifold, ignoring noise or redundant dimensions. Previous work has found that neural networks have low ID, capturing data in low-dimensional manifolds [48, 60]. Following [22], for a given threshold $\beta$, the intrinsic dimension is the $d\in\mathbb{N}$ such that the ratio of explained variance for $d$ dimensions of a $N$ dimensional PCA is above $\beta$:
$\frac{\sum_{i=1}^{d}var(y_{i})}{\sum_{j=1}^{N}var(y_{j})}>\beta$ (9)
First, we measure the ID of both unguided and guided Deep FCNs across all layers at different points of training; see fig.˜14. Crucially, at 300 steps of training, we notice that guidance with an untrained guide network preserves the initial ID found in the randomly initialized Deep FCN representations. The trained guide network and base training achieve low ID values, which is consistent with findings in prior work [11]. At the end of training, all networks reach the same intrinsic dimension. This finding indicates that the dynamics of training, as shown by ID, change with or without guidance. The Deep FCN without guidance achieves a low ID too early in training, and this is likely similar for the guidance with a trained guide network.
One interpretation of these results is that ID has an effect on overfitting in the Deep FCN. When ID is too low, the Deep FCN overfits. Therefore, we can discover a new regularization scheme for training the Deep FCN based on ID. In this scheme we introduce a new loss function, which designs a differentiable version of PCA-ID based on a specific ID threshold and forces the ID of the representations to be above a particular ID. In particular, given the activations from a specific layer of the target network, ${\bm{A}}^{T}_{i}$, a variance threshold $\tau$, and a target ID $t$, we first find the SVD of the activations,
${\bm{A}}^{T}_{i}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}$ (10)
We extract the eigenvalues using the singular values, $\lambda_{j}=\Sigma_{jj}^{2}$ and find the total variance using the eigenvalues.
$T=\sum_{j=1}^{r}\lambda_{j}+\epsilon$ (11)
where $r$ is the total number of nonzero singular values and $\epsilon$ is a small constant for numerical stability. Afterwards, we find the explained ratios and cumulative sums of the eigenvalues:
$p_{j}=\frac{\lambda_{j}}{T}~,~c_{k}=\sum_{j=1}^{k}p_{j}$ (12)
We use the cumulative sum to find the loss for being below the target ID in eq.˜13.
$\ell_{\text{below}}=\sum_{j=1}^{t-1}\sigma(\beta(c_{j}-\tau))$ (13)
where $\sigma$ is the sigmoid function and $\beta$ controls the sharpness of the loss. Similarly, eq.˜14 gives the loss for being too far above the target ID.
$\ell_{\text{above}}=\sigma(\beta(\tau-c_{t}))$ (14)
Figure 15: Controlling intrinsic dimensionality of Deep FCN representations improves image classification performance. We introduce a novel loss function to regularize the ID of Deep FCN representations in each linear layer of the network during training. We find that this leads to improved validation accuracy, similar to guidance. This shows that guidance correlates with geometric modifications to target networks and can find new regularizations.
Our total loss is given by $\mathcal{L}=\ell_{\text{above}}+\ell_{\text{below}}$. Using this loss to control ID, we tune each layer in our Deep FCN and see a final validation loss given in fig.˜15. We find that using our new loss function to control the ID of every linear layer of our Deep FCN leads to improved validation performance and an accuracy of 14.65%. This aligns with guidance, indicating that guidance may be controlling the ID of the target network. Furthermore, we have used guidance to find a new regularization scheme based on intrinsic dimensionality.
I.3 Linear Decoding
Figure 16: ImageNet classes are barely decodable from a randomly initialized ResNet-18. In order to assess the performance of our randomly initialized networks, we design a linear decoder to decode ImageNet classes from all layers of the network. Chance accuracy is 0.1% (1/1000) on the graph above. We find that, while we can decode ImageNet classes with an accuracy above chance, the accuracy is still very low.
We assess whether object classes are decodable from internal representations of a randomly initialized ResNet-18. A potential explanation for improvements in the target network with an untrained guide is the ability to linearly decode the ImageNet classes with significant accuracy at certain layers of the untrained guide network.
We train a linear decoder with 4000 ImageNet images from the train set and test on 1000 images from the validation set for each layer. We show results in fig.˜16. We find that the linear decoder never achieves any accuracy greater than 0.8% for any of the layers. Furthermore, later layers contain little information that is useful to linearly decode ImageNet classes. This means that linearly decodable information isn’t present in the guide network and this aspect isn’t driving improvements in target networks. We note that this matches findings in Amid et al. [2], which reports that linear decodability from a ResNet-18 achieves 3.4% top-1 accuracy. The increase in performance is likely due to using a much larger dataset.
Appendix J Guidance with RSA and Ridge Regression
J.1 Representational Similarity Analysis
@@ I.3 Linear Decoding
We use the RSA formulation as described in Kriegeskorte et al. [46]. Specifically, RSA constructs representational dissimilarity matrices (RDMs) for two sets of representations and compares them using an outer similarity function.
@@
As with CKA, we use the complement of the similarity to construct $\bar{\mathcal{M}}$. In practice, we define $d$ to be the cosine distance between every pair of inputs and $s$ to be the pearson correlation between the RDMs as done in previous work [14].
@@
We apply guidance with RSA to Deep FCN as our target network and ResNet-18 as our guide network. Similar to our CKA results, we train for 100 epochs with a batch size of 256, as RSA is sensitive to the number of samples when comparing sets of representations.
J.2 Ridge Regression
We used similar ridge regression formulation as [14, 71] without cross-validation.
Given two sets of representations, ${\bm{R}}\in\mathbb{R}^{b\times d_{1}}$ and ${\bm{R}}^{\prime}\in\mathbb{R}^{b\times d_{2}}$, we first apply a sparse random projection on the representations. Since the dimensionality of the representations is prohibitively large, the projection makes the ridge regression feasible to compute. We refer to the resulting representations as $\bm{P}$} and $\bm{P}$$^{\prime}$ which correspond to the projected representations $\bm{R}$} and $\bm{R}$$^{\prime}$ respectively. $\bm{P}$} and $\bm{P}$} have dimension $d$, where $d$ is fixed using the Johnson-Lindenstrauss lemma [40].
Afterwards, we mean-center the representations and apply ridge regressions using the original least-squared solution as follows. Our goal is to predict the representations $\bm{P}$$^{\prime}$ using regressors over $\bm{P}$}. We first split our representations into a training set i.e. ${\bm{P}}_{\text{train}},{\bm{P}}^{\prime}_{\text{train}}$ and testing set ${\bm{P}}_{\text{test}},{\bm{P}}^{\prime}_{\text{test}}$ where the training set contain half the representations and the testing set contains the other half. We first a set of regressors $\hat{\beta}$ as follows:
$\hat{\beta}=(({\bm{P}}_{\text{train}})^{T}{\bm{P}}_{\text{train}}+\lambda{\bm{I}}_{d})^{-1}({\bm{P}}_{\text{train}})^{T}{\bm{P}}^{\prime}_{\text{train}}$ (17)
where $\lambda$ is the ridge penalty, which is a hyperparameter. The coefficients $\hat{\beta}$ are then used to predict the held out data where:
$\begin{gathered}\hat{{\bm{P}}^{\prime}_{\text{test}}}={\bm{P}}_{\text{test}}\beta\end{gathered}$ (18)
We measure the cosine similarity between the predicted representations $\hat{{\bm{P}}^{\prime}_{\text{test}}}$ and actual representations ${\bm{P}}_{\text{test}}$.
$\mathcal{M}({\bm{R}},{\bm{R}}^{\prime})=\text{cosine}(\hat{{\bm{P}}^{\prime}_{\text{test}}},{\bm{P}}^{\prime}_{\text{test}})$ (19)
We apply guidance with RSA to Deep FCN as our target network and ResNet-18 as our guide network. Similar to our CKA results, we train for 100 epochs with a batch size of 256, as ridge regression is sensitive to the number of samples when comparing sets of representations. We manually tune the $\lambda$ hyperparameter, finding that $\lambda=10.0$ is optimal for the trained guide network and $\lambda=100.0$ is optimal for the untrained guide network.
J.3 Results
Figure 17: Guidance with RSA as the representational similarity metric maintains similar performance to CKA. We include a further experiment where we change the metric for representational alignment from CKA to RSA during guidance training. We apply this to the Deep FCN with ResNet-18 as a guide network. We see that, like CKA, RSA alignment also allows for transferring the prior from ResNet-18. However, unlike CKA, the untrained guide network only does marginally better than the trained network, potentially indicating the RSA is better at transferring trained features.
Figure 18: Guidance with ridge regression as the representational similarity metric improves performance over CKA. We change the metric for representational alignment from CKA to ridge regression during guidance training. We apply this to the Deep FCN with ResNet-18 as a guide network. We see that, like CKA, ridge regression alignment also allows for transferring the prior from ResNet-18. We find that this improves over CKA significantly
RSA: We see results over the training, validation, and representational dissimilarity loss in fig.˜17. The Deep FCN guided by a trained ResNet-18 achieves an accuracy of 11.02% and the Deep FCN guided by a randomly initialized ResNet-18 achieves an accuracy of 11.74%.
@@
We can first observe that guided training improves over base training as noted in fig.˜2 and table˜3. This demonstrates the generality of our approach to other metrics. As long as a representational similarity metric is differentiable, we can optimize the metric for alignment between two networks as a method to transfer the prior of one network to another.